R=100 XL = 200 E= 100V L 30° Xc = 1302 I What is ZT in polar form? 10 0 25 11.18 Q 226.6 50 210 11.18 0 263.43
Answer #1

R=lor XL=202 M Xo=152. I E=100/30 (1 27 = R + JX2 - } Xo 27 = 10 + 8 (20-15) = (10+55).. 127) = J 102+52 5125 :)).182 127 tan